Prev Next

SW&IT Meeting Minutes 2020-06-01

Attendees: Carl, David, Igor, Matthieu, Petr, Stephen, Xianyu, Yang

Outlook

Labor Reporting

Report auto-generated every Monday at 6:30am, Do not edit this table, Update your entries in smartsheets.

20200518
LR20200518 - Bec, Matthieu 73.0
LR20200518 - Holmberg, Carl 40.0
LR20200518 - Edmeades, David 40.0
LR20200518 - Sola, Igor 72.0
LR20200518 - Hooper, Stephen 80.0
LR20200518 - Zhang, Yang 40.0
LR20200518 - Zhang, Xianyu 40.0

News from management

Previous Action Items

Owner Action Due DateSorted ascending Completion Date
IS return macbook-pro to Steward or DE 20200601  

Open topics

  • AGw3 recovery (email from Thomas Hahn)
    • vpn/ssh(?), mountain inspection, wiki update with Thomas Hahn's instructions to upload new f/w
  • 4x screens in Al's office
  • s/w weekly back to 10:30am

Team reports

PK

MB

  • Last week
    • OC, Management
    • SHARK-VIS updated RFA continued
    • DMS 'deep' refactoring and cleaning
    • ROWG screen tests + AnyDesk
    • TMS meeting
  • This week
    • DMS
      • websockets, AdSec, etc.
      • LBC stiched png preview
      • max_bytes for ovms
      • merge IMS into DMS
    • cobbler / kickstart, ims f30
    • RPM/CI short tutorial

DE

  • In possession of MBP and spare Luci webio device
    • Ordered adapter to use Dell power supply to power webio
    • Will configure if needed
  • Removed Wade from telescope work email list after consulting with him and Pat. I am assured that he is supplied with important emails.
  • Pat is having trouble accessing IMS from VA. He is VPN'd in, other internal resources work, but IMS times out. I verified at the same time that it works for me. Is there some kind of communication timeout that is so short that the roundtrip across the country is outside the range?
  • Bucki proceeding with Ketiv contract renewal
  • New safety training
IT topics:
  • Jim managed to install his printer!
  • AnyDesk runs in the background; let's make sure that users fully exit from program
  • Finally convinced Jim to authenticate with his UA account for O365. His nag issue is now resolved
  • Still nothing from Crucial; will follow up
  • Waiting for CDW re: Veeam quote
  • Luci server spec ongoing
  • Cisco training in progress
  • Hearing reports that HD manufacturers are slipping inappropriate technology (SMR) into "NAS" rated drive lines. Will see how this develops and if any of our drives are affected. Gives credence to the otherwise frustrating "approved drive" lists.
  • Avtech progress?
  • Engineering student Hannah software help
  • obs2 / space increase

XZ

  • Last week
    • Troubleshooting of SX ASM issue
    • AdSec restart
    • Meetings
  • This week
    • Write reset shell procedure with Guido
    • Make oaa_lib backwards compatible for new4D
    • Look more data for the SX ASM issue
    • PB action items

YZ

  • Previous:
    • Some AO-Cravat work (Steve Ertel has LBTO github access).
    • TMS wiki startup and meeting (needs to determine the TMS software meeting schedule).
    • Investigated Olga's MODS UI request (need to rewrite some MODS script, obsN machines seem to have "xdpyinfo" installed, so seems doable).
    • Misc things (DMS MODS checking, MODS alarm script restore on obs3 etc.)
    • Safety training.
  • Next:
    • LBTO wiki access for the TMS team members?
    • AO-Cravat meeting Friday.
    • TMS software meeting (expand wiki material)?
    • Try to make new scripts for Olga's MODS UI request.
    • Arrange with Olga to test LBC operation from 64-bit machine.
    • Continue previous Python projects.

SH

  • set up for obs in Tucson
  • Powered on an plumbed "new" VPN server
  • worked on RPMS
  • Google LDAP to replace AD
    • shell pushed up (seems to work)
    • attempts with Synology not so succesful, but I got users from LDAP to show up. There is an upgrade that would allow going directly to Google rather than going through proxy
    • FreeRADIUS looks good.
    • OpenLDAP account merge
    • UID matching on abell and oingo
  • Laptops to Leroy (also the desktop)
  • AGW3
    • call with Christian B., Dan R., Dan C. Ilya, James
    • Ilya wanting firmware
  • Business drive changed access
  • Olga and robs32 -- told her to forget it, but most of her problems were related to "lbto" being bash and not csh. I rewrote the cshrc to be bashrc.
  • meetings

next:

  • satellite image change
  • more Google LDAP playing
    • VPN

IS

Last week:
  • LUCI
    • work on #IT7952
    • work on configuration of LUCI WebIO
  • PIT
    • If I find some time, work on the PIT admin tool backend.
  • OT
    • Move QA backend to pit.lbto.org and update release pages.
    • work on MODs backend issues
    • Carl helped me out understanding his RPM and CI server configuration files.

This week:
  • Safety Training
  • LUCI
    • work on #IT7952
    • work on configuration of LUCI WebIO, the HTTP GET api seems to be the same but on this device seems to be not consistent. Have to work a bit more on it. David E. provided support for network installation and backup configuration if needed.
    • Need to partner with Jenny and revert/test to old GEIRS in LUCI2 osa-luci.sim
  • PIT
    • If I find some time, work on the PIT admin tool backend.
  • OT
    • Probably more backend work related to MODS XSLT

CH

Last Week:
  • DMS: reorganized new sensor handling code for branch.
  • MODS: assisted Olga with finding azcam images, was looking on mods-cam vnc desktops rather than azcam.
  • OA Control: documented PMAC code surrounding axis encoder, email to John Morris, Dan Cox, PK for feedback.

This Week:
  • AGW: email encoder handing query to Omron, get encoder/board order status from JM.
  • AHS: restart work on LBTI alarm patch.
  • Argos: get RPM build working.
  • DMS: commit new sensor handling code to branch.
  • MODS: check if there's a more elegant way to check MODS AGW and guide camera connectivity after restarts.
  • OA Control: email Omron/Delta Tau regarding expected value from binary vs gray encoder.
  • Safety training
Topic revision: r19 - 03 Jun 2020, CarlHolmberg
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback